### v3

# ZooplanktonBench: A Geo-Aware Zooplankton Recognition and Classification Dataset from Marine Observations

Abstract.
Plankton are small drifting organisms found throughout the world’s oceans and can be indicators of ocean health. One component of this plankton community is the zooplankton, which includes gelatinous animals and crustaceans (e.g. shrimp), as well as the early life stages (i.e., eggs and larvae) of many commercially important fishes. Being able to monitor zooplankton abundances accurately and understand how populations change in relation to ocean conditions is invaluable to marine science research, with important implications for future marine seafood productivity. While new imaging technologies generate massive amounts of video data of zooplankton, analyzing them using general-purpose computer vision tools turns out to be highly challenging due to the high similarity in appearance between the zooplankton and its background (e.g., marine snow). In this work, we present the ZooplanktonBench, a benchmark dataset containing images and videos of zooplankton associated with rich geospatial metadata (e.g., geographic coordinates, depth, etc.) in various water ecosystems. ZooplanktonBench defines a collection of tasks to detect, classify, and track zooplankton in challenging settings, including highly cluttered environments, living vs non-living classification, objects with similar shapes, and relatively small objects.Our dataset presents unique challenges and opportunities for state-of-the-art computer vision systems to evolve and improve visual understanding in dynamic environments characterized by significant variation and the need for geo-awareness. The code and settings described in this paper can be found on our website: https://lfk118.github.io/ZooplanktonBench_Webpage.
Computer vision, Object detection, Fine-grained classification, Video analysis, Marine science
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737395isbn: 979-8-4007-1454-2/2025/08submissionid: dtbfb0064ccs: Computing methodologies Object recognitionccs: Computing methodologies Object detectionccs: Computing methodologies Trackingccs: Applied computing Environmental sciences
1. Introduction
Plankton are small drifting organisms found throughout the world’s oceans . One component of this plankton community is the zooplankton, which includes gelatinous animals and crustaceans (e.g. shrimp), as well as the young stages of many commercially important fishes. Because zooplankton are a critical component of marine food webs, there is a strong scientific motivation to monitor their abundances accurately and understand how their abundances change in relation to ocean conditions which is important for future seafood productivity prediction.
Imaging technologies are used to measure zooplankton abundances in situ with high spatial resolution, yet these technologies have also introduced many challenges (Benfield et al., 2007; Lombard et al., 2019). In productive shallow water ecosystems (<200 m depth, where most commercial and recreational fishing occurs), zooplankton often live within dense patches of detrital material, also known as “marine snow” due to its fluffy or stringy appearance and gradual sinking speed. This marine snow includes dead and decaying phytoplankton (plankton that performs photosynthesis similar to terrestrial green plants), zooplankton fecal material, organic mucus, or other excretions from organisms (Turner, 2015). In situ imaging technologies have demonstrated that marine snow is highly patchy and can also have a variety of appearances making it difficult to differentiate zooplankton from marine snow. In addition, zooplankton have a variety of morphologies. The young fishes and other larval animals are especially uncommon yet critical to identify accurately within the sea of marine snow common to biologically productive and economically important habitats.
These characteristics set up a challenge for machine vision in the ocean. How can we accurately separate living organisms from the much more common and variable marine snow? Can we accurately identify rarer organisms like larval fishes and more common groups simultaneously? Sometimes organisms within the images are interacting, and detecting those interactions automatically may be important for understanding zooplankton ecology (Greer et al., 2021). Accomplishing these goals would greatly enhance the ability to measure and understand ocean plankton populations. Solving these problems would have other applications for large imaging datasets where targets exist in environments with variable noise. Creative approaches are needed to combine available datasets (human annotated) and videos of live animals to accomplish these tasks.
In this paper, we propose ZooplanktonBench dataset that features a rich image and video dataset of zooplankton associated with geospatial metadata in various water ecosystems. Challenging benchmark tasks and tracks are designed to recognize zooplankton in complex environments. Our dataset presents unique challenges and opportunities for state-of-the-art computer vision systems. We expect our data can be beneficial to both the marine science and machine learning communities.
2. Related Work
Efforts to automatically identify plankton using machine vision have been underway for over 30 years (Hu and Davis, 2005; Sosik and Olson, 2007), with more recent efforts gravitating towards the use of Convolutional Neural Networks (CNNs) (Luo et al., 2018; Schmid et al., 2023). Ellen et al. (Ellen et al., 2019) found that CNNs performed better when the plankton images were coupled with “context metadata,” which includes information from oceanographic sensors (i.e., the depth, temperature, salinity, and other ocean variables associated with each image). These results seem reasonable given that marine animals are patchy (Robinson et al., 2021; Greer et al., 2016), meaning that one type is detected by an imaging system, and similar types are nearby that are also similar in size. Ideally, classification should rely solely on the image itself, just as an ocean researcher looks at specimens to make identifications (without bias as to where the organism came from). Nevertheless, the environmental origin of the training data could be important.
3. ZooplanktonBench Data
Our ZooplanktonBench dataset is collected from the northern Gulf of Mexico (nGOM), which is a highly biologically productive system that fuels fisheries and coastal economies in the United States. One reason for that high biological productivity is the high amount of river discharge that enters into the nGOM carrying high amounts of nutrients. These nutrients fuel the growth of photosynthesizing phytoplankton, which serve as the food for some types of zooplankton, and that energy eventually makes it up to productive fisheries (Grimes, 2001). The earliest stages of fishes are known as larvae, and they are 1 mm to several cm in size and are considered members of the zooplankton community. This study took place underneath a river plume, so the images from the shallowest depths (10 m) are just underneath the plume and feature large amounts of dead and decaying phytoplankton, known as marine snow. At the deeper depths, 25m and 35m, the images have less marine snow because much of the snow has been consumed or broken down.
Our data was collected with a towed plankton imaging vehicle, the In Situ Ichthyoplankton Imaging System (see Figure 1(a)), on July 24, 2011 (Cowen and Guigand, 2008). It was towed behind a ship at 3 fixed depths in nGOM just east of the Louisiana Bird Foot Delta (Greer et al., 2016) (see Figure 2). The ISIIS uses a small point light source, two plano-convex lenses, and a line-scan camera to collect images at 16Hz (Cowen and Guigand, 2008). The vehicle is towed at a constant speed of 2.5 meters per second and easily maintain a constant depth (as was done in this study). The images are associated with oceanographic sensor data, such as geographic coordinates, depth, temperature, salinity, dissolved oxygen, and chlorophyll-a fluorescence (which serves as a proxy for phytoplankton abundance). These datasets provide a high-resolution map of oceanographic properties and how the organisms respond to these conditions. The images are collected with the bottom two pods that house a camera and a light source. The marine creatures and marine snow passing between the pods block the light and are photographed as a “shadow”. Images were transferred via a fiber optic cable to a shipboard computer where they were stored on hard drives. The system has real-time information on the vehicle position, the associated oceanographic sensor data, and live-streaming data. Figure 1(b) and 1(c) show an example of an eel larva (Leptocephalus larva) captured in a container as well as its image view in a shadowgraph imaging system.
(a) ISIIS
(b) Eel larva
(c) Eel larva image
Figure 1. (a) The In Situ Ichthyoplankton Imaging System (ISIIS, left) on the deck of a research vessel. Plankton or particles that pass in between the bottom two pods are captured in the images as "shadows." (b) An example of an eel larva (Leptocephalus larva) captured in a container. (c) The same eel larva as viewed in a shadowgraph imaging system.
Figure 2. The region of sampling in the nGOM. Bathymetry contours correspond to 10-m intervals (20, 30, 40, and 50 m contours are labeled). The red line corresponds to the star-shaped sampling pattern between the 40 and 50-m isobaths, including crossings at three different depths (10, 25, and 35 m). The bottom right inset provides a 3D view of the sampling region.
[…]
(a) Chaetognath
(b) Chaetognath
(c) Hydromedusa
(d) Hydromedusa
(e) Lobate ctenophore
(f) Lobate ctenophore
(g) Shrimp
(h) Shrimp
(i) Unknown
Figure 3. Example photos of zooplankton in our ZooplanktonBench dataset.
Table 1. Detailed annotation information of ZooplanktonBench.
Zooplankton Class Number of annotations Annotation from 10 meters Annotation from 25 meters Annotation from 35 meters
Chaetognath or Arrow worm 235,065 66,874 49,030 119,161
Shrimp 17,243 7,656 7,286 2,301
Fish larva 2,432 1,011 772 649
Scyphomedusa 856 571 269 16
Pleurobrachia 823 606 138 79
Hydromedusa or Jellyfish 7,821 2,033 983 4,805
Thaliacea 452 81 279 91
Stomatopod larva 1,024 474 502 48
Lobate ctenophore 671 343 153 175
Siphonophore 2994 774 948 1,272
Polychaete worm 852 14 66 772
Ctenophore or Comb jelly 22 10 10 2
Pteropod 82 3 68 11
Cumacean 2,418 0 0 2418
Crab zooea or Crab larva 1,336 0 0 1,336
Copepod 7,366 0 356 7,010
Unknown 4,276 884 1,703 1,689
Total 285,733 81,334 62,564 141,835
In addition to the towed imaging dataset, videos were collected of organisms in some categories onboard a ship using a benchtop imaging system (Greer et al., 2025). Live animals that we captured in a net and placed into a small container onboard a research vessel in preparation for filming. The videos feature zooplankton swimming around in a container at a frame rate of 25 Hz. The idea behind these videos was that they would be useful for quantifying swimming behaviors and building robust training libraries. The latter goal has yet to be evaluated but could be useful for augmenting training libraries and automatically detecting living animals in images (and separating them from marine snow).
This high-quality human-annotated dataset features a few other characteristics that make it valuable for computer vision. The contrasting environments with different marine snow amounts are representative of a typical biologically productive environment. So identifying the marine animals in these variable environments is an important task for understanding the connection of marine animals to these habitats, and the task is also representative of a general problem in computer vision in diverse and cluttered backgrounds. The identification categories are highly imbalanced (chaetognaths are very common, for example) and some categories, such as fish larvae, have many different morphologies because they belong to different species or stages of development. Together, the features of this dataset represent an opportunity to employ the latest techniques for object detection and recognition in variable backgrounds. On a practical level for ocean scientists, it is useful for living organisms to be separated from the marine snow, thus giving scientists a simpler dataset to annotate. If video is useful for improving image classification, that would be a major step forward for efficiency.
4. ZooplanktonBench Tasks and Tracks
The imagery data collected in Section 3 presents rich geo-aware information and great challenges to computer vision systems. We design several tasks for ZooplanktonBench, which focus on the unique characteristics of the data. For each of the learning tasks, we make multiple tracks to train and evaluate computer vision models under different scenarios and setups. In Section 5, we show the performance of state-of-the-art CV systems on our data and point out their limitations and challenges for computer vision and machine learning. Notably, the ZooplanktonBench dataset contains rich information that can be used for other computer vision tasks such as self-supervised video learning, or object tracking in complex scenes. We expect our published dataset can be found useful by those researchers as well.
4.1. Tasks
Our primary objective is to detect and classify zooplankton given imagery data collected in Section 3. We design the following learning-based tasks that are not only invaluable for marine science research but also challenging for the deep learning community.
4.1.1. Fine-grained detection and classification on zooplankton species
As discussed in Section 1, being able to correctly detect and classify zooplankton from images is invaluable to marine science. However, the appearance of the different zooplankton types is highly variable, making it a challenging task for learning systems. In this task, we provide fine-grained zooplankton labels and their bounding boxes for each image and ask an object detection system to locate and classify them in new unseen images. This fine-grained classification task is a fundamental problem in image classification and object detection in computer vision. We adopt a similar training setup and widely used evaluation metrics such as Mean Average Precision (50, 50-95) to measure the performance of various methods.
Figure 4. Examples of zooplankton and marine snow. Marine snow is abundant and can take many different shapes. By chance, they can resemble living zooplankton in the images.
4.1.2. Living zooplankton vs. marine snow detection
One of the unique challenges in our data/task is the predominance existence of distractors (i.e., non-living marine snow particles) in the data. As shown in Figure 4, there are a lot of objects that are visually similar to the zooplankton but they are actually belonging to “marine snow.” The high similarity in their appearances highlights the difficulty of differentiating living zooplankton from marine snow. The marine snow consists of biological debris and organic material falling from upper waters to the deep ocean (Oceanic and Administration, 2020). To make matters more challenging, in the video data, not only zooplankton can move in the water, the marine snow objects can also move in motion due to the disturbance of the water around them.
[…]
4.2. Tracks
4.2.1. Diverse environments
As described in Section 3, the samples we collected are from different depths in the ocean: 10 meters, 25 meters, and 35 meters. As a result, the characteristics of the data change significantly across the depths. Even though zooplankton appearance remains consistent across these depths, the imaging quality, environmental complexity, and variations pose great challenges to the detection and classification task. Figure 5 shows example figures at the various depths.
(a) 10 meters
(b) 25 meters
(c) 35 meters
Figure 5. Image examples from 10 meters, 25 meters, and 35 meters. Each image is 13 cm by 13 cm and looks through 40 cm of ocean water (i.e., image depth of field).
In this track setup, we evaluate the performance of a computer vision system on the tasks in Section 4.1 in diverse environments with differences in the background abundance of targets and marine snow. Specifically, we measure the classification and detection performance of a CV model that is trained on {10 meters, 25 meters, 35 meters, all-mix} datasets and tested on {10 meters, 25 meters, 35 meters, and all-mix} datasets. This setup generates a comprehensive view of the accuracy and generalization ability of a model trained on the ZooplanktonBench data, enabling an evaluation of its geo-aware performance with respect to ocean depth.
4.2.2. Images + video
Otherwise purely relaying on images (denoted as "Image Only"), it has been shown in prior research that training with a mix of fully supervised and unsupervised data can boost the performance of machine learning systems (Berthelot et al., 2019). In particular, training with additional video data can help downstream single-image tasks (Parthasarathy et al., 2023). In our case, we augment the fully labeled image training set with the unlabeled video data.
In this track, we evaluate the performance of a model on the tasks in Section 4.1 with the training of both the fully supervised image data and unlabeled video data on classification and detection. To simplify the preparation of the training setup, we provide an efficient object tracking tool, ByteTrack (Zhang et al., 2022), to work closely with any trained object detector on our data. In Section 5, we describe a basic pipeline to leverage video data in the training framework of a fine-tuned object detector.
5. Experiments
Here we show the performance of representative state-of-the-art detection and classification computer vision algorithms trained and tested on our data with the tasks and tracks defined in Section 4. For each method, we also discuss their limitations and challenges. All the experiments are conducted on a Linux Ubuntu workstation equipped with an NVIDIA RTX A6000 GPU with 48GB. The average training time for one fine-tuned model is 15 hours.
5.1. Fine-tuned object detector
YOLO (You Only Look Once) is a cutting-edge object detection model (Redmon et al., 2016) that has evolved over multiple iterations and gained widespread adoption in various applications due to its speed and accuracy (Lou et al., 2023; Talaat and ZainEldin, 2023; Xiao et al., 2024). In this work, we select the recently released YOLOv8 (Jocher et al., 2023) as the base object detector. We fine-tune YOLOv8 for two tasks: fine-grained detection/classification and living zooplankton vs. marine snow detection.
For the image + video track, we integrated the fine-tuned YOLOv8 model with ByteTrack to produce object tracking results. These results will include labeled objects, with some objects assigned an “id” and others without one. The “label” is provided by YOLOv8, while the “id” is assigned by ByteTrack. An object assigned an “id” indicates that it is trackable and consistently detected as the same class across multiple frames (5 frames in our experiment). This consistency suggests a higher likelihood of the object being a living zooplankton compared to objects without it. Based on this assumption, we can enhance the detection accuracy by disregarding labeled objects that are untracked.
Results. Table 2 shows YOLOv8’s results on the fine-grained zooplankton detection and classification task under the diverse environments track. Table 3 shows YOLOv8’s results on the living zooplankton vs. marine snow detection task under the regular image-only and the images + video track. Figure 6 shows examples of detection results of living zooplankton and fine-grained classification results of zooplankton using the trained model.
Figure 6. YOLOv8 detection and classification results. Top: living zooplankton detection. Bottom: fine-grained classification.
Table 2. Experimental results of YOLOv8 on the fine-grained zooplankton species classification task under the diverse environment track. “mix” refers to the combined dataset from all three depths—10 meters, 25 meters, and 35 meters.
Training dataset Testing dataset mAP50 mAP50-95
10 meters 10 meters 0.303 0.162
10 meters 25 meters 0.328 0.164
10 meters 35 meters 0.128 0.0547
10 meters mix data 0.193 0.0991
25 meters 10 meters 0.182 0.0941
25 meters 25 meters 0.594 0.354
25 meters 35 meters 0.179 0.0813
25 meters mix data 0.229 0.124
35 meters 10 meters 0.0799 0.0389
35 meters 25 meters 0.12 0.0738
35 meters 35 meters 0.411 0.284
35 meters mix data 0.138 0.0839
mix data 10 meters 0.328 0.173
mix data 25 meters 0.619 0.358
mix data 35 meters 0.486 0.252
mix data mix data 0.434 0.228
Table 3. Experimental results of YOLOv8 on the living zooplankton vs. marine snow detection task.
Track Model mAP50 mAP50-95
Image Only YOLOv8 0.871 0.522
Images + video YOLOv8 + ByteTrack 0.84 0.475
Discussion. Table 2 shows that, except for the model trained at 35 meters, all other models achieve the best results on the 25-meter dataset. As previously discussed in Section 4.2, environmental complexity varies with depth. Images at the 25-meter depth are characterized by an environment that is relatively sparse (compared to 10 meters) and may have zooplankton composition most similar to the other 2 depths. In addition, 25 meters appears to be the most balanced environment for the YOLOv8 detector. Further examination of performance on the “mix” dataset reveals that the model trained on the 25 meters dataset outperforms those trained at 10 and 35 meters in mAP, thereby reaffirming the importance of the 25 meters data. Together, the model trained on the mix depth data significantly outperforms the model trained solely at 25 meters across all test cases, suggesting that integrating data from various depths enhances the model’s accuracy and generalizability. A more effective model could be developed by carefully mixing data from different depths.
Challenges. Fine-tuning an object detector on custom data is one of the most widely used computer vision applications. While the YOLOv8 detectors show reasonable performance in fine-grained classification and detection, the absolute numbers in mAP50-95 are significantly worse than that of the general object detection results (such as those in the COCO benchmark (Lin et al., 2014)). It is known that common general object detectors could suffer from objects with small scale and in cluttered backgrounds (Hoiem et al., 2012; Liu et al., 2020). Therefore we expect the challenges exist in our tasks and tracks can help the community to design better object detection algorithms that can benefit the general object tasks.
5.2. Open-set object detector
Recently, open-set object detectors such as Grounding DINO (Liu et al., 2023b) have gained popularity largely due to their flexibility. Once trained, these detectors can output a detection result with an arbitrary object prompt that is potentially unseen during training. This contrasts with traditional classifiers that operate under the closed-set assumption—namely, that the test sample’s ground truth label will belong to one of the categories pre-defined during the training phase. We adopt the Grounding DINO model and evaluate its open-set detection performance on ZooplanktonBench in a zero-shot manner.
Results. We found that the confidence threshold for the open-set detectors is a critical factor in the final performance. This threshold must be adjusted based on the dataset, implying that a single threshold may not be universally effective across different scenarios. In our context, a single threshold setting proved inadequate for varying depths and categories within the marine environment.
Table 4 shows the quantitative results of the Grounding DINO’s zero-shot performance on ZooplanktonBench’s zooplankton fine-grained classification task with different confidence thresholds. Table 5 shows the zero-shot performance of the Grounding DINO on ZooplanktonBench’s zooplankton detection task under different confidence thresholds.
Table 4. The zero-shot performance of Grounding DINO on the zooplankton fine-grained classification task under different confidence thresholds.
Depth Threshold 0.1 Threshold 0.075 Threshold 0.05
mAP50 mAP50-95 mAP50 mAP50-95 mAP50 mAP50-95
10 meters 0.2 0.0429 0.1809 0.0382 0.1229 0.0257
25 meters 0.0978 0.0212 0.0852 0.0189 0.0439 0.01
35 meters 0.0159 0.0051 0.0125 0.0039 0.0082 0.0023
Table 5. The zero-shot performance of Grounding DINO on living zooplankton detection task under different confidence thresholds.
Depth Threshold 0.1 Threshold 0.075 Threshold 0.05
mAP50 mAP50-95 mAP50 mAP50-95 mAP50 mAP50-95
10 meters 0.1352 0.0312 0.1106 0.0251 0.0585 0.0138
25 meters 0.0628 0.0123 0.0208 0.009 0.0267 0.0053
35 meters 0.0274 0.0057 0.046 0.0045 0.0125 0.0027
Discussion. Table 4 shows that a higher threshold generally correlates with improved performance, indicating a positive relationship between the threshold and the performance. Moreover, an examination of Table 5 reveals that performance at 25 meters does not consistently decline with a reduced threshold. Although the overall trend aligns with that observed in Table 4, caution is still necessary. This reveals how sensitive such open-set detectors can be with the proper setup of the thresholds. Sometimes this could be a challenge to end users. Besides the confidence threshold, the performance of Grounding DINO on different depth datasets from ZooplanktonBench varies significantly. For instance, the mAP50 scores at 10-meter, 25-meter, and 35-meter datasets, with a threshold of 0.1, are 0.2, 0.0978, and 0.0159 respectively. The mAP decreases by 51.1% when moving from 10 meters to 25 meters and by 83.7% from 25 meters to 35 meters. Unlike YOLOv8 which shows the best performance at 25 meters, Grounding DINO is more effective at 10 meters. This shows the different characteristics of the object detectors.
Challenges. Besides the sensitivity of the threshold discussed above, one major challenge we found about the open-set detector is the prompt-based detection design. To detect unseen inputs, the model incorporates a language decoder that assists in interpreting these inputs by categorizing them into known classes. Although this feature minimizes the need for extensive pre-training, it introduces challenges in accurate classification. Specifically, in our dataset, the detector struggles to distinguish between “larval fish,” “shrimp,” and “chaetognath,” possibly because the decoder erroneously groups these distinct inputs into similar or identical classes. As a result, the classification performance is considerably compromised. For example, as illustrated in Table 4, the model’s performance metrics are disappointing. The mAP50 is only 0.2 and the mAP50-95 is 0.0429 on 10-meter data with a confidence threshold of 0.1. In contrast, the YOLOv8 model, trained specifically for our dataset, significantly outperforms the open-set detector, achieving a mAP50 of 0.328 and a mAP50-95 of 0.173 on the same test dataset. We hope our data and benchmark can be used as a testbed to evaluate the generalization ability of future open-set object detectors.
5.3. Large visual-language model
Inspired by the success of foundational models such as GPT (Radford et al., 2019), large visual-language models have been developed and demonstrated impressive results on many visual reasoning tasks (Liu et al., 2023a). We evaluate the performance of GPT-4V from Microsoft Azure OpenAI (OpenAI, 2024) on our geo-aware dataset ZooplanktonBench. Unlike traditional approaches with YOLO or Grounding DINO, where the original images are directly processed for detection, we employ a different methodology with GPT-4V: we first run the object detector on the input image and crop it according to the detected box. This strategy allows GPT-4V to perform our fine-grained zooplankton classification tasks, thereby enhancing the robustness of the results from the object detectors without significant additional computational costs.
We design two prompts to the GPT4-V model, the first prompt is as follows: “You are a Marine Biologist. I am sending you an image containing zooplankton. The possible classes are chaetognath, larval fish, hydromedusa, lobate ctenophore, pleurobrachia, shrimp, siphonophore, stomatopod larva, unknown, thaliac, polychaete worm, scyphomedusa, other ctenophore. Please give me the most probable name shown by the given image in the following format: class name. Please just give a short answer and do not provide explanations for your choice.”
We also design a second prompt to explicitly encode geo-metadata to evaluate the performance of GPT-4V being geo-aware (Mac Aodha et al., 2019; Mai et al., 2020; Yang et al., 2022; Mai et al., 2023a, b; Cole et al., 2023; Wu et al., 2024): “You are a Marine Biologist. I am sending you an image containing zooplankton. The image was captured at latitude 29.5063167 and longitude 88.5215167, situated 34.5 meters beneath the seafloor. The environmental conditions at this location were as follows: an irradiance of 1.70E-05, a water pressure of 34.73 dbar, a temperature of 22.21 degrees Celsius, a salinity of 36.36 psu, and a dissolved oxygen level of 1.8678 mg per L. Given these conditions, the possible classes of zooplankton are: chaetognath, larval fish, hydromedusa, lobate ctenophore, pleurobrachia, shrimp, siphonophore, stomatopod larva, unknown, thaliac, polychaete worm, scyphomedusa, other ctenophore. Please give me the most probable name shown by the given image in the following format: class name. Please just give a short answer and do not provide explanations for your choice.“
The underlined part will be replaced with the corresponding geo-metadata.
The detected-and-cropped images are passed directly to GPT-4V immediately following the prompt. GPT-4V then classifies each image crop by responding with one of the predefined classes or stating "I can’t assist you" if it cannot recognize the image. Any response that does not match the established categories is classified as an error message.
Table 6. GPT-4V fine-grained classification results on ZooplanktonBench. ‘Correct’ indicates the number of images correctly classified. ‘Classification’ indicates the number of images predicted by GPT-4V, the superscript on ‘Correct’ and ‘Classification’ indicates the data sent to GPT-4V, and ‘Instances’ shows the total number of images in each class in ground truth data.
Class CorrectImage ClassificationImage CorrectImage+Geo ClassificationImage+Geo Instances
Chaetognath 587 1533 423 1194 666
Larval fish 78 114 86 137 432
Hydromedusa 36 109 9 45 162
Lobate ctenophore 11 118 14 146 18
Pleurobrachia 0 0 0 0 45
Shrimp 3 29 3 22 362
Siphonophore 9 37 2 10 108
Stomatopod larva 11 13 6 10 126
Unknown 0 6 0 15 21
Thaliac 0 0 0 0 17
Polychaete worm 1 108 3 254 6
Scyphomedusa 19 20 20 20 216
Other ctenophore 3 17 0 12 9
Results. We selected a dataset of 2,188 representative zooplankton images from ZooplanktonBench for GPT-4V. Out of the total, without the input of geo-metadata, GPT-4V successfully classified 758 images but encountered 84 error messages during the classification process. With the input of geo-metadata, GPT-4V successfully classified 566 images but encounter 323 error messages during the classification process. The results of this classification are detailed in Table 6.
Discussion. For the fine-grained classification task, GPT-4V seems to struggle with the concept of zooplankton. Unlike its impressive results for classifying common objects such as cats and dogs, classifying zooplankton requires domain knowledge in marine science. Our ZooplanktonBench could be a great supplement to such a need. With the zooplankton images alone, GPT-4V does performance well. In Table 6, “Chaetognath” is identified in 70.1% of the total images, which is nearly triple the actual number of real chaetognath images present. This discrepancy suggests that GPT-4V struggles to accurately distinguish chaetognath from other types of zooplankton. This misclassification might be attributed to the widespread distribution of chaetognaths relative to other classes, increasing their likelihood of being represented in the training dataset utilized by GPT-4V. Additionally, an examination of the ‘Correct’ and ‘Instances’ columns reveals a high ratio of correctly identified chaetognath images to their total instances. This suggests that while GPT-4V frequently misclassifies other zooplankton as chaetognath, it successfully recognizes a substantial number of actual chaetognath instances.
As mentioned earlier, incorporating geo-metadata into GPT-4V increases error messages from 84 to 323, suggesting that additional information may not necessarily improve the model’s performance. As shown in Table 6, compared to GPT-4V using images alone, geo-metadata improves accuracy for “Lobate ctenophore” by 16.67% and “Larval fish” by 1.85% but reduces accuracy for “Chaetognath” by 24.62%, despite fewer false positives. However, false detections increase for “Lobate ctenophore” by 25 and “Larval fish” by 15. “Polychaete worm” is heavily over-detected, with false positives rising from 108 to 254. “Pleurobrachia” and “Thaliac” remain undetected, and the “Unknown” category shows more detections without improved accuracy. Overall, geo-metadata enhances classification in some cases but reduces accuracy for “Chaetognath” and “Hydromedusa”.
Challenges. While it has often been discussed that large foundational models are “running out of data” to be trained on (out, 2022), our evaluation shows that general-trained large models can struggle on data that require domain knowledge and fine-grained classification. We hope ZooplanktonBench can be used to evaluate the generalization ability of large visual and language models, maybe with a knowledge base type of approach: for example, by resorting to the literature in marine science, the model can learn the appearance features of different zooplankton types.
6. Conclusion
We present ZooplanktonBench, a carefully designed and curated geo-aware dataset that features challenging and unique computer vision tasks such as detection and fine-grained classification of marine science imagery data. Our data features zooplankton that are highly diverse in appearance, and their habitat can be replete with with somewhat similar-looking marine snow particles. Detection and classification systems developed in ZooplanktonBench will be useful in the marine science community and may have applications for advancing computer vision and deep learning systems.
Several tests and developments were not explored in this paper that may have value in terms of scientific advancements and societal relevance. One limitation of ZooplanktonBench is that we do not provide data annotation for our benchtop video dataset. Future work of ZooplanktonBench may involve using annotated still images to develop an AI-assisted video annotation tool to help with this process. Improving algorithms to address the computer vision challenges illustrated in this benchmark dataset will help with the automated monitoring of marine populations and developing action plans to protect these ecologically and economically valuable areas of the planet.
Acknowledgements
This work was supported by funds from the University of Georgia Presidential Interdisciplinary Seed Grant. Gengchen Mai acknowledges support from the Microsoft Research Accelerate Foundation Models Academic Research (AFMR) Initiative. Adam Greer is supported by funding from the U.S. National Science Foundation (OCE-2244690 and OCE-2023133).